### v1

@@ 5.4 Learning Stability Analysis
Figure 5 reports the coefficient of variation (CoV = std/mean) over the last 200 training epochs for the value-function loss and the estimated advantages. Lower is better.
@@ 5.5 Generalization to Unseen Conditions
We train in the Thin Obstacle environment and evaluate zero-shot on more challenging, previously unseen terrains. Table 6 reports means and standard deviations over seeds. For vision-only policies, collision counts are omitted because the agents move too little for this metric to be informative.
Table 6: Performance on challenging unseen terrains (zero-shot from Thin Obstacle).
Model Architecture Reward Collisions Distance Moved
@@
Transformer Vision-Only 151.75 $\pm$ 82.28 - 5.75 $\pm$ 2.68
Mamba Vision-Only 28.53 $\pm$ 29.25 - 2.79 $\pm$ 0.96
@@
Cross-modal policies generalize substantially better than single-modality baselines. Relative to Proprio-Only, LocoMamba increases return by approximately $448\%$ (from $106.45$ to $583.76$), reduces collisions by $39.7\%$ (from $779.40$ to $470.27$), and raises progress distance by $386.6\%$ (from $4.94$ to $24.04$). Compared with Transformer Proprio-Vision, LocoMamba achieves a $126.8\%$ higher return, $20.8\%$ fewer collisions, and a $46.1\%$ longer distance. The vision-only variants yield low displacement, indicating weak generalization without proprioceptive anchoring. These out-of-distribution results show that combining proprioception with depth is effective for generalization and that the Mamba SSM improves cross-modal fusion under distribution shift.
@@ 6 Conclusion
Across challenging simulated environments, LocoMamba achieved higher returns, higher success rates, fewer collisions, and longer progress distances than strong baselines that include proprioception-only controllers, depth-only controllers, recurrent fusion, and Transformer-based fusion. Learning-curve analyses and efficiency metrics showed faster convergence and improved sample efficiency under the same compute budget, consistent with the near-linear scaling of the selective state-space backbone. Stability indicators, including variability of advantages and value loss, further confirmed reliable optimization.
Ablation studies established the importance of combining proprioception with depth and showed that the Mamba backbone is particularly effective for cross-modal fusion. The proposed PPO training protocol, which combines domain randomization, curriculum scheduling, and a compact reward, produced stable learning and balanced trade-offs among progress, smoothness, and safety. Together, these results answer the research questions on overall performance, efficiency, modality design, and training robustness in the affirmative.
Due to current budgetary and hardware-access constraints, we have not conducted real-world experiments. When resources permit, we will deploy LocoMamba on a quadruped platform and assess sim-to-real transfer, latency, and safety under field conditions.

### v2

@@ 5.4 Learning Stability Analysis
Fig. 5 reports the coefficient of variation (CoV = std/mean) over the last 200 training epochs for the value-function loss and the estimated advantages. Lower is better.
@@ 5.5 Generalization to Unseen Conditions
The policies are trained in the Thin Obstacle environment and evaluate them zero-shot on two previously unseen scenarios: a rugged terrain and a dynamic-obstacle terrain with randomly moving hazards. Tables 6 and 7 report means and standard deviations over 10 runs. For vision-only policies, collision counts are omitted because the agents move too little for this metric to be informative.
Table 6: Zero-shot performance on a rugged unseen terrain (trained on Thin Obstacle). Mean $\pm$ std over 10 runs.
Model Architecture Return Collision Times Distance Moved
@@
Transformer Vision-Only 151.75 $\pm$ 82.28 5.75 $\pm$ 2.68
Mamba Vision-Only 28.53 $\pm$ 29.25 2.79 $\pm$ 0.96
@@
Table 7: Zero-shot performance on a dynamic-obstacle terrain (trained on Thin Obstacle). Mean $\pm$ std over 10 runs.
Model Architecture Return Collision Times Distance Moved
Proprio-Only 119.20 $\pm$ 26.33 423.03 $\pm$ 187.08 5.01 $\pm$ 0.88
Transformer Vision-Only 196.78 $\pm$ 44.64 7.49 $\pm$ 1.51
Mamba Vision-Only 19.32 $\pm$ 26.52 2.57 $\pm$ 0.79
Transformer Proprio-Vision 685.49 $\pm$ 133.31 121.63 $\pm$ 83.15 28.36 $\pm$ 4.77
LocoMamba (Ours) 749.09 $\pm$ 269.19 38.47 $\pm$ 43.96 32.51 $\pm$ 5.84
Rugged Terrain.
Cross-modal policies generalize substantially better than single-modality baselines. LocoMamba achieves the highest return and progress with the fewest collisions. Relative to Proprio-Only, return increases by approximately $448\%$, collisions decrease by $39.7\%$, and distance increases by $386.6\%$. Compared with Transformer Proprio-Vision, LocoMamba attains a $126.8\%$ higher return, $20.8\%$ fewer collisions, and a $46.1\%$ longer distance. Vision-only variants travel little, indicating limited generalization without proprioceptive anchoring. These results show that combining proprioception with depth is effective for out-of-distribution generalization and that the selective state-space backbone strengthens cross-modal fusion under distribution shift.
Dynamic obstacles.
LocoMamba maintains a clear advantage when dynamically moving. Relative to Transformer Proprio-Vision, return increases by about $9\%$, collisions are reduced by roughly two thirds, and distance increases by about $15\%$. Relative to Proprio-Only, return and distance increase by more than a factor of six while collisions fall by approximately $91\%$. Vision-only policies again show limited displacement.
These findings indicate that the proposed PPO-trained cross-modal Mamba design preserves performance under temporal disturbances and moving hazards, demonstrating strong generalization to previously unseen challenging and dynamic environments. Taken together, the results address RQ1 to RQ4 by confirming overall performance gains, the effectiveness of combining proprioception and depth, the efficiency of the Mamba state-space backbone, and the robustness of the training protocol.
@@ 6 Conclusion
Across challenging simulated environments, LocoMamba achieved higher returns, fewer collisions, and longer progress distances than strong baselines that include proprioception-only controllers, depth-only controllers, and Transformer-based fusion. Learning-curve analyses and efficiency metrics showed faster convergence and improved sample efficiency under the same compute budget, consistent with the near-linear scaling of the selective state-space backbone. Stability indicators, including variability of advantages and value loss, further confirmed reliable optimization.
Ablation studies established the importance of combining proprioception with depth and showed that the Mamba backbone is particularly effective for cross-modal fusion. The proposed PPO training protocol, which combines domain randomization, curriculum scheduling, and a compact reward, produced stable learning and balanced trade-offs among task-aligned progress, energy efficiency, and safety. Together, these results answer the research questions on overall performance, efficiency, modality design, and training robustness in the affirmative.
Due to current budgetary and hardware-access constraints, the real-world experiments have not been conducted. When resources permit, LocoMamba will be deployed on a quadruped platform and assess sim-to-real transfer, latency, and safety under field conditions.